# Adaptive LoRA Experts Allocation and Selection for Federated Fine-Tuning

### v1

@@ 3.3 Observations
As shown in fig.˜2, our experimental results indicate that in Scenario 1, FedIT performs significantly worse than FedLoRA-MoE, suggesting that clients with different tasks struggle to contribute effectively to a single shared LoRA module. This observation is consistent with findings in centralized settings [14, 37], where a single LoRA module proves insufficient for handling diverse domains. Conversely, in Scenario 2, FedIT outperforms FedLoRA-MoE, suggesting that using separate LoRA experts for homogeneous clients may be redundant and can degrade performance while increasing inference overhead, which aligns with results from [8]. These findings offer key insights into expert allocation strategies under varying task heterogeneity in FL.
@@
We observe that task similarity between clients can be effectively assessed by computing the cosine similarity of their LoRA $B$ matrices after a short period of local training. To validate this, we conducted an experiment with four clients: two using the SST-2 dataset and two using the QNLI dataset. As shown in fig.˜3, clients working on the same task develop highly similar LoRA $B$ matrices, while those working on different tasks exhibit significantly lower similarity. Interestingly, this pattern is exclusive to the LoRA $B$ matrices; the $A$ matrices show no consistent relationship with task similarity. This observation supports findings in [37, 16], which suggest that the output transformation matrix $B$ captures task-specific information, whereas the input matrix $A$ tends to encode general linguistic features shared across tasks. Although a similar task-specific pattern can be observed by analyzing the product $BA$, this requires matrix multiplication and full-rank projection recovery, which incurs significantly higher computational cost. In contrast, using the $B$ matrices alone provides a lightweight yet effective proxy for task similarity, making it more practical in FL.
@@
To investigate our second research question, we conducted additional experiments under the task-heterogeneous setting (Scenario 1) using the FedLoRA-MoE approach. We varied the top-$k$ parameter in the MoE router, testing values of $k=2,3,4$ (where $k=4$ corresponds to using all available experts). As shown in fig.˜4, different clients achieve optimal performance at different $k$ values—some benefit most from $k=2$, while others perform better with $k=3$ or $k=4$. This result highlights that even if the total number of trained LoRA experts is fixed, clients have varying needs regarding how many experts they should utilize. A static top-$k$ selection strategy is therefore suboptimal across all clients. These findings motivate the design of an adaptive expert selection mechanism that dynamically determines the optimal number of experts for each client based on its specific data characteristics.
@@ 4 Proposed Method
In this section, we describe the framework design of our proposed method FedLEASE by explaining how it addresses the two important challenges in complex federated fine-tuning settings. A theoretical analysis of the proposed method can be found in appendix˜E.
@@ 4.1 Adaptive LoRA Experts Allocation
We evaluate the quality of each clustering configuration using the silhouette coefficient [34], which measures how well each client fits within its assigned cluster relative to others. The average silhouette score for a $k$-cluster configuration is defined as $S(k)=\frac{1}{N}\sum_{i=1}^{N}s^{k}(i)$, where $s^{k}(i)$ is the silhouette coefficient for client $i$ given by $s^{k}(i)=\frac{b^{k}(i)-a^{k}(i)}{\max(a^{k}(i),b^{k}(i))}$. Here, $a^{k}(i)$ denotes the average dissimilarity between client $i$ and all other clients in the same cluster (intra-cluster cohesion), while $b^{k}(i)$ is the minimum average dissimilarity between client $i$ and clients in other clusters (inter-cluster separation). The optimal number of experts $M=\underset{2\leq k\leq M_{max}}{\arg\max}S(k)$ is selected as the $k$ that maximizes the average score. This approach ensures that we identify both the optimal number of experts needed and the most coherent grouping of clients based on their adaptation patterns. A high silhouette score indicates that clients within the same cluster exhibit similar adaptation characteristics, while clients in different clusters show distinct patterns. Once the optimal clustering $\mathcal{C}^{M}=\{C^{M}_{1},C^{M}_{2},\ldots,C^{M}_{M}\}$ is determined, we initialize each LoRA by aggregating the clients within the corresponding cluster:
@@ 4.2 Adaptive top-$M$ LoRA Experts
An illustration of the adaptive top-$M$ mechanism is shown in fig.˜5. This design allows each client to dynamically select between 1 and $M$ distinct experts based on input characteristics. When the router favors positions [0,M1][0,M{-}1], the client primarily uses its assigned expert; when favoring [M,2M1][M,2M{-}1], it leverages multiple or all experts. Intermediate combinations—e.g., selecting $p$ copies of the assigned expert and MpM{-}p others—result in using Mp+1M{-}p{+}1 unique experts. This mechanism ensures all clients follow a uniform top-$M$ protocol, yet adaptively vary the number of experts used. The assigned expert is always included, enabling effective updates. Thus, our approach eliminates manual tuning and supports data-driven expert selection tailored to each client.
@@ 4.3 Algorithm Workflow
FedLEASE operates in two phases: initialization and iterative training. A detailed algorithm˜1 can be found in appendix˜B.
@@ 5.1 Natural Language Understanding
We evaluate our method on four GLUE [40] benchmark datasets: SST-2, QNLI, MRPC, and QQP. Unlike prior works such as FedSA, which assume clients differ only in label distribution, we adopt a more realistic heterogeneous setting, where each client is assigned a different dataset from the four. The setup includes 16 clients in total, with four clients per dataset. Each client’s data is randomly partitioned from the full corresponding dataset. We use RoBERTa-Large (355M) [30] from the HuggingFace library as the base model. Additional training details are provided in appendix˜A.
@@
Ablation on Adaptive Expert Allocation. FedLEASE assigns 16 clients to 4 experts based on clustering results (illustrated in appendix˜C). To evaluate the impact of expert training allocation, we compare against the following alternatives: FedLoRA-Single: A single expert is trained with contributions from all clients. To ensure fair comparison, we test two variants with different LoRA ranks: FedLoRA-Single ($r=4$) and FedLoRA-Single ($r=16$). FedLoRA-Individual: Each client trains a separate expert, resulting in 16 experts and one-to-one client-expert mapping. FedLEASE (w/o adaptive top-$M$): Uses the same expert allocation as FedLEASE but replaces adaptive top-$M$ router with vanilla fixed top-2. Results in table˜2 show that our clustering-based expert allocation achieves the best performance, even without the adaptive top-$M$ mechanism. Both FedLoRA-Single variants underperform due to limited capacity to model heterogeneous data, while FedLoRA-Individual, despite higher computational cost, still lags behind FedLEASE, validating the effectiveness of our clustering strategy in balancing knowledge sharing and task specificity.
We also evaluate router aggregation strategies. Compared to maintaining individual routers per client, our approach—averaging router networks within each group—achieves better performance (fig.˜6(a)), confirming the benefit of shared routing among clients with similar data.
@@
Ablation on Adaptive top-$M$ Mechanism. We assess the effectiveness of our adaptive top-$M$ mechanism by comparing it with fixed top-$k$ strategies (top-1 through top-4). As shown in fig.˜6(b), different clients achieve optimal performance with different $k$ values—for example, clients with QQP dataset perform best with top-2, whereas those with MRPC dataset benefit more from top-4. Notably, all fixed top-$k$ strategies except top-$1$ achieve comparable performance and outperform baseline methods reported in table˜1, also highlighting the importance of the expert allocation strategy. In contrast, our adaptive top-$M$ mechanism consistently surpasses all fixed strategies across clients, demonstrating its capability to dynamically select the optimal number of experts per input. To further illustrate this adaptivity, fig.˜6(c) visualizes expert selection across layers for 16 clients in our main experiment. The patterns show substantial variation both across clients and across layers within the same client—some layers rely on a single expert, while others combine multiple. Client ID 0-3, 4-7, 8-11 and 12-15 correspond to SST-2, QNLI, MRPC and QQP, respectively. Specifically, we can observe that: 1. Vertical trend: As layers deepen, the average number of activated LoRA experts tends to increase; 2. Horizontal trend: As task difficulty increases (reflected by decreasing accuracy from SST-2 to QQP in table˜1), the average number of activated LoRA experts also increases; 3. Cluster-level: Since we aggregate the router within each cluster group during server aggregation, rather than maintaining each client’s individual one, clients within the same cluster tend to exhibit similar but not identical expert selection patterns. This fine-grained adaptivity underscores the limitations of fixed top-$k$ approaches and confirms the necessity of our adaptive top-$M$ mechanism.
@@ 5.2 Natural Language Generation
In addition to NLU tasks, we evaluate our method on natural language generation (NLG) tasks. We adopt LLaMA-2-7B [39] from Hugging Face as the base model and use four FLAN datasets—Text Editing, Struct to Text, Sentiment Analysis, and Commonsense Reasoning—to construct a heterogeneous client setting. We consider 8 clients in total, with each dataset assigned to two clients. Training details are provided in appendix˜A. As shown in table˜3, FedLEASE consistently outperforms all baselines on NLG tasks. Compared to the strongest overall baseline, FedLEASE achieves gains of 2.26%, 0.46%, 1.43%, and 1.85% on Text Editing, Struct to Text, Sentiment Analysis, and Commonsense Reasoning, respectively. On average, FedLEASE improves by 1.50%, highlighting its ability to handle heterogeneous client data across both classification and generation tasks, demonstrating its generalizability beyond NLU.
@@ 6 Conclusion
Appendix A Training Details
NLU Task. We consider 16 clients in total, with four clients assigned to each of the four GLUE datasets. Each client’s data is randomly partitioned from the corresponding full dataset. RoBERTa-Large (355M) [30] (24 transformer layers) from HuggingFace is used as the base model. AdamW is adopted as the optimizer for all methods, with a batch size of 128, local epochs set to 2, and a total of 25 communication rounds. Following [35], LoRA is applied to the query and value projections in the attention layers, and the classification head is frozen after initialization. For our method, the upper bound of experts $M_{max}$ is set to 8 and the LoRA rank to 4. Baselines are configured to ensure comparable computational workloads. Learning rates are selected via grid search from $\eta\in\{1\text{E}{-4},3\text{E}{-4},5\text{E}{-4},1\text{E}{-3},3\text{E}{-3},5\text{E}{-3}\}$. Accuracy is utilized as the evaluation metric for NLU tasks.
NLG Tasks. For NLG tasks, we use LLaMA-2-7B [39] with 8-bit quantization (32 transformer layers) from Hugging Face as the base model and select four FLAN datasets—Text Editing, Struct to Text, Sentiment Analysis, and Commonsense Reasoning—to construct a heterogeneous client setting. A total of 8 clients are considered, with each dataset assigned to two clients. Each client has 600 training samples and 200 test samples. All methods use AdamW as the optimizer, with a batch size of 8, local epochs set to 2, 10 communication rounds and the upper bound of experts $M_{max}$ is set to 8. LoRA is applied to the query and value matrices in the attention layers, with a LoRA rank of 8. Learning rates are selected via grid search from $\eta\in\{1\text{E}{-4},3\text{E}{-4},1\text{E}{-3},3\text{E}{-3},1\text{E}{-2}\}$. Followed by [44], we choose ROUGE-1 as the evaluation metric for NLG tasks.
Appendix B The Algorithm of FedLEASE
@@
Appendix C More Experiment Results
C.1 Clustering Results
Our clustering analysis reveals natural groupings of clients based on the cosine similarity of their LoRA B matrices. Figure 7 presents a comprehensive visualization of the clustering results using three complementary approaches.
Figure 7(a) shows the Silhouette scores for different numbers of clusters, ranging from $k=2$ to $k=8$. The Silhouette score, which measures how similar objects are to their assigned cluster compared to other clusters, peaks at $k=4$, indicating that 4 is the optimal number of clusters for the main experiments setting. This finding suggests that clients naturally form 4 distinct groups based on their model adaptations. In Figure 7(b), we visualize the distance matrix derived from cosine similarity between client LoRA B matrices. The heatmap reveals clear block diagonal structures, which further supports the existence of distinct client clusters. The darker squares along the diagonal represent groups of clients with high intra-cluster similarity, while lighter colors indicate greater dissimilarity between different clusters.
Finally, Figure 7(c) presents a hierarchical clustering dendrogram based on the same cosine similarity measure. The dendrogram provides an alternative view of client relationships, illustrating how clients progressively merge into larger groups. The vertical axis represents the distance at which clusters are combined, with longer vertical lines indicating greater separation between clusters.

### v2

@@ 3.3 Observations
As shown in Figure˜2, our experimental results indicate that in Scenario 1, FedIT performs significantly worse than FedLoRA-MoE, suggesting that clients with different tasks struggle to contribute effectively to a single shared LoRA module. This observation is consistent with findings in centralized settings [14, 37], where a single LoRA module proves insufficient for handling diverse domains. Conversely, in Scenario 2, FedIT outperforms FedLoRA-MoE, suggesting that using separate LoRA experts for homogeneous clients may be redundant and can degrade performance while increasing inference overhead, which aligns with results from [8]. These findings offer key insights into expert allocation strategies under varying task heterogeneity in FL.
@@
We observe that task similarity between clients can be effectively assessed by computing the cosine similarity of their LoRA $B$ matrices after a short period of local training. To validate this, we conducted an experiment with four clients: two using the SST-2 dataset and two using the QNLI dataset. As shown in Figure˜3, clients working on the same task develop highly similar LoRA $B$ matrices, while those working on different tasks exhibit significantly lower similarity. Interestingly, this pattern is exclusive to the LoRA $B$ matrices; the $A$ matrices show no consistent relationship with task similarity. This observation supports findings in [37, 16], which suggest that the output transformation matrix $B$ captures task-specific information, whereas the input matrix $A$ tends to encode general linguistic features shared across tasks. Although a similar task-specific pattern can be observed by analyzing the product $BA$, this requires matrix multiplication and full-rank projection recovery, which incurs significantly higher computational cost. In contrast, using the $B$ matrices alone provides a lightweight yet effective proxy for task similarity, making it more practical in FL.
@@
To investigate our second research question, we conducted additional experiments under the task-heterogeneous setting (Scenario 1) using the FedLoRA-MoE approach. We varied the top-$k$ parameter in the MoE router, testing values of $k=2,3,4$ (where $k=4$ corresponds to using all available experts). As shown in Figure˜4, different clients achieve optimal performance at different $k$ values—some benefit most from $k=2$, while others perform better with $k=3$ or $k=4$. This result highlights that even if the total number of trained LoRA experts is fixed, clients have varying needs regarding how many experts they should utilize. A static top-$k$ selection strategy is therefore suboptimal across all clients. These findings motivate the design of an adaptive expert selection mechanism that dynamically determines the optimal number of experts for each client based on its specific data characteristics.
@@ 4 Proposed Method
In this section, we describe the framework design of our proposed method FedLEASE by explaining how it addresses the two important challenges in complex federated fine-tuning settings. A theoretical analysis of the proposed method can be found in Appendix˜F.
@@ 4.1 Adaptive LoRA Experts Allocation
We evaluate the quality of each clustering configuration using the silhouette coefficient [34], which measures how well each client fits within its assigned cluster relative to others. The average silhouette score for a $k$-cluster configuration is defined as $S(k)=\frac{1}{N}\sum_{i=1}^{N}s^{k}(i)$, where $s^{k}(i)$ is the silhouette coefficient for client $i$ given by $s^{k}(i)=\frac{b^{k}(i)-a^{k}(i)}{\max(a^{k}(i),b^{k}(i))}$. Here, $a^{k}(i)$ denotes the average dissimilarity between client $i$ and all other clients in the same cluster (intra-cluster cohesion), while $b^{k}(i)$ is the minimum average dissimilarity between client $i$ and clients in other clusters (inter-cluster separation).
The optimal number of experts $M=\underset{2\leq k\leq M_{max}}{\arg\max}S(k)$ is selected as the $k$ that maximizes the average score. This approach ensures that we identify both the optimal number of experts needed and the most coherent grouping of clients based on their adaptation patterns. A high silhouette score indicates that clients within the same cluster exhibit similar adaptation characteristics, while clients in different clusters show distinct patterns. Once the optimal clustering $\mathcal{C}^{M}=\{C^{M}_{1},C^{M}_{2},\ldots,C^{M}_{M}\}$ is determined, we initialize each LoRA by aggregating the clients within the corresponding cluster:
@@ 4.2 Adaptive top-M LoRA Experts
An illustration of the proposed adaptive top-$M$ mechanism is shown in Figure˜5. The proposed router $\hat{\omega}\in\mathbb{R}^{2M-1}$ allows each client to decide, for every input, how many and which experts contribute, instead of relying on a globally fixed $k$. When the top-ranked scores lie among the first $M$ entries of $\hat{\omega}$, the computation is dominated by the client’s own assigned expert $E_{j}$. When large scores appear in the remaining M1M{-}1 positions, the client leverages additional experts. Intermediate cases arise naturally: if the router selects $p$ of the first $M$ entries (i.e., $p$ internal components of the assigned expert $E_{j}$) together with MpM{-}p entries from the other experts, then in effect Mp+1M{-}p{+}1 unique experts participate in the forward computation.
To make the mechanism more concrete, consider a case with M=3M{=}3 experts $\{E_{1},E_{2},E_{3}\}$ and a client whose assigned expert is $E_{1}$. The router produces
$\hat{\omega}=\big[\underbrace{\hat{\omega}_{1}^{E_{1}},\,\hat{\omega}_{2}^{E_{1}},\,\hat{\omega}_{3}^{E_{1}}}_{\text{connected to assigned expert }E_{1}},\;\underbrace{\hat{\omega}_{4}^{E_{2}},\,\hat{\omega}_{5}^{E_{3}}}_{\text{connected to other experts}E_{2},E_{3}}\big]$
where the first three entries are independently learned routing scores corresponding to distinct internal components of the assigned expert $E_{1}$ (all routed to $B_{1}A_{1}x$), and the last two correspond to the other experts $E_{2}$ and $E_{3}$ (routed to $B_{2}A_{2}x$ and $B_{3}A_{3}x$). With different input samples inducing different routing score distributions, our proposed top-$M$ mechanism can accordingly select different expert combinations:
  • Sample 1: if the top-3 are $\hat{\omega}_{1}^{E_{1}},\hat{\omega}_{2}^{E_{1}},\hat{\omega}_{3}^{E_{1}}$, then only the assigned expert $E_{1}$ contributes.
  • Sample 2: if the top-3 are $\hat{\omega}_{1}^{E_{1}},\hat{\omega}_{2}^{E_{1}},\hat{\omega}_{4}^{E_{2}}$, then two unique experts $\{E_{1},E_{2}\}$ are selected.
  • Sample 3: if the top-3 are $\hat{\omega}_{2}^{E_{1}},\hat{\omega}_{4}^{E_{2}},\hat{\omega}_{5}^{E_{3}}$, then all three experts $\{E_{1},E_{2},E_{3}\}$ participate.
This mechanism guarantees that each client’s designated expert $E_{j}$ always participates—facilitating stable local updates—while allowing flexible, data-driven cooperation with other experts. By enabling the router to balance the internal components of $E_{j}$ with the contributions from the remaining $\{E_{m}\}_{m\neq j}$, the method adapts to both input complexity and inter-client heterogeneity, avoiding any manual tuning of $k$ and achieving effective expert utilization for federated fine-tuning.
@@ 4.3 Algorithm Workflow
FedLEASE operates in two phases: initialization and iterative training. A detailed Algorithm˜1 can be found in Appendix˜A.
@@ 5 Experiment
5.1 Training Details
NLU Task. We consider 16 clients in total, with four clients assigned to each of the four GLUE datasets. Each client’s data is randomly partitioned from the corresponding full dataset. RoBERTa-Large (355M) [30] (24 transformer layers) from HuggingFace is used as the base model. AdamW is adopted as the optimizer for all methods, with a batch size of 128, local epochs set to 2, and a total of 25 communication rounds. Following [35], LoRA is applied to the query and value projections in the attention layers, and the classification head is frozen after initialization. For our method, the upper bound of experts $M_{max}$ is set to 8 and the LoRA rank to 4. Baselines are configured to ensure comparable computational workloads. Learning rates are selected via grid search from $\eta\in\{1\text{E}{-4},3\text{E}{-4},5\text{E}{-4},1\text{E}{-3},3\text{E}{-3},5\text{E}{-3}\}$. Accuracy is utilized as the evaluation metric.
NLG Tasks. For NLG tasks, we use LLaMA-2-7B [39] with 8-bit quantization (32 transformer layers) from Hugging Face as the base model and select four FLAN datasets—Text Editing, Struct to Text, Sentiment Analysis, and Commonsense Reasoning—to construct a heterogeneous client setting. A total of 8 clients are considered, with each dataset assigned to two clients. Each client has 600 training samples and 200 test samples. All methods use AdamW as the optimizer, with a batch size of 8, local epochs set to 2, 10 communication rounds and the upper bound of experts $M_{max}$ is set to 8. LoRA is applied to the query and value matrices in the attention layers, with a LoRA rank of 8. Learning rates are selected via grid search from $\eta\in\{1\text{E}{-4},3\text{E}{-4},1\text{E}{-3},3\text{E}{-3},1\text{E}{-2}\}$. Followed by [44], we choose ROUGE-1 as the evaluation metric.
@@ 5.2 Natural Language Understanding
We evaluate our method on four GLUE [40] benchmark datasets: SST-2, QNLI, MRPC, and QQP. Unlike prior works such as FedSA, which assume clients differ only in label distribution, we adopt a more realistic heterogeneous setting, where each client is assigned a different dataset from the four. The setup includes 16 clients in total, with four clients per dataset. Each client’s data is randomly partitioned from the full corresponding dataset. We use RoBERTa-Large (355M) [30] from the HuggingFace library as the base model. Additional training details are provided in Section˜5.1.
@@
Ablation on Adaptive Expert Allocation. FedLEASE assigns 16 clients to 4 experts based on clustering results (illustrated in Appendix˜B). To evaluate the impact of expert training allocation, we compare against the following alternatives: FedLoRA-Single: A single expert is trained with contributions from all clients. To ensure fair comparison, we test two variants with different LoRA ranks: FedLoRA-Single ($r=4$) and FedLoRA-Single ($r=16$). FedLoRA-Individual: Each client trains a separate expert, resulting in 16 experts and one-to-one client-expert mapping. FedLEASE (w/o adaptive top-$M$): Uses the same expert allocation as FedLEASE but replaces adaptive top-$M$ router with vanilla fixed top-2. Results in Table˜2 show that our clustering-based expert allocation achieves the best performance, even without the adaptive top-$M$ mechanism. Both FedLoRA-Single variants underperform due to limited capacity to model heterogeneous data, while FedLoRA-Individual, despite higher computational cost, still lags behind FedLEASE, validating the effectiveness of our clustering strategy in balancing knowledge sharing and task specificity.
We also evaluate router aggregation strategies. Compared to maintaining individual routers per client, our approach—averaging router networks within each group—achieves better performance (Figure˜6(a)), confirming the benefit of shared routing among clients with similar data.
@@
Ablation on Adaptive top-$M$ Mechanism. We assess the effectiveness of our adaptive top-$M$ mechanism by comparing it with fixed top-$k$ strategies (top-1 through top-4). As shown in Figure˜6(b), different clients achieve optimal performance with different $k$ values—for example, clients with QQP dataset perform best with top-2, whereas those with MRPC dataset benefit more from top-4. Notably, all fixed top-$k$ strategies except top-$1$ achieve comparable performance and outperform baseline methods reported in Table˜1, also highlighting the importance of the expert allocation strategy. In contrast, our adaptive top-$M$ mechanism consistently surpasses all fixed strategies across clients, demonstrating its capability to dynamically select the optimal number of experts per input. To further illustrate this adaptivity, Figure˜6(c) visualizes expert selection across layers for 16 clients in our main experiment. The patterns show substantial variation both across clients and across layers within the same client—some layers rely on a single expert, while others combine multiple. Client ID 0-3, 4-7, 8-11 and 12-15 correspond to SST-2, QNLI, MRPC and QQP, respectively. Specifically, we can observe that: 1. Vertical trend: As layers deepen, the average number of activated LoRA experts tends to increase; 2. Horizontal trend: As task difficulty increases (reflected by decreasing accuracy from SST-2 to QQP in Table˜1), the average number of activated LoRA experts also increases; 3. Cluster-level: Since we aggregate the router within each cluster group during server aggregation, rather than maintaining each client’s individual one, clients within the same cluster tend to exhibit similar but not identical expert selection patterns. This fine-grained adaptivity underscores the limitations of fixed top-$k$ approaches and confirms the necessity of our adaptive top-$M$ mechanism.
@@ 5.3 Natural Language Generation
In addition to NLU tasks, we evaluate our method on NLG tasks. We adopt LLaMA-2-7B [39] as the base model and use four FLAN datasets—Text Editing, Struct to Text, Sentiment Analysis, and Commonsense Reasoning—to construct a heterogeneous client setting. We consider 8 clients in total, with each dataset assigned to two clients. Training details are provided in Section˜5.1. As shown in Table˜3, FedLEASE consistently outperforms all baselines on NLG tasks. Compared to the strongest overall baseline, FedLEASE achieves gains of 2.26%, 0.46%, 1.43%, and 1.85% on Text Editing, Struct to Text, Sentiment Analysis, and Commonsense Reasoning, respectively. On average, FedLEASE improves by 1.50%, highlighting its ability to handle heterogeneous client data across both classification and generation tasks, demonstrating its generalizability beyond NLU.
@@ 6 Conclusion
Acknowledgments
The work of Lei Wang, Jieming Bian and Jie Xu is partially supported by NSF under grants 2433886, 2505381 and 2515982. The work of Letian Zhang is partially supported by NSF under grant 2348279 and also supported by MTSU Stark Land project.
Appendix A The Algorithm of FedLEASE
@@
Appendix B Clustering Results
Our clustering analysis reveals natural groupings of clients based on the cosine similarity of their LoRA B matrices. Figure˜7 presents a comprehensive visualization of the clustering results using three complementary approaches.
Figure˜7(a) shows the Silhouette scores for different numbers of clusters, ranging from $k=2$ to $k=8$. The Silhouette score, which measures how similar objects are to their assigned cluster compared to other clusters, peaks at $k=4$, indicating that 4 is the optimal number of clusters for the main experiments setting. This finding suggests that clients naturally form 4 distinct groups based on their model adaptations. In Figure˜7(b), we visualize the distance matrix derived from cosine similarity between client LoRA B matrices. The heatmap reveals clear block diagonal structures, which further supports the existence of distinct client clusters. The darker squares along the diagonal represent groups of clients with high intra-cluster similarity, while lighter colors indicate greater dissimilarity between different clusters.
Finally, Figure˜7(c) presents a hierarchical clustering dendrogram based on the same cosine similarity measure. The dendrogram provides an alternative view of client relationships, illustrating how clients progressively merge into larger groups. The vertical axis represents the distance at which clusters are combined, with longer vertical lines indicating greater separation between clusters.